$SPAD/src/input mathml.input

Arthur Ralfs

December 3, 2016

Abstract

Contents

— * —

)set break resume
)spool mathml.output
)set message test on
)set message auto off
)clear all
--S 1 of 21
(x+y)**2
--R
--R
--R         2           2
--R   (1)  y  + 2x y + x
--R                                                    Type: Polynomial(Integer)
--E 1
--S 2 of 21
coerce(%)$MMLFORM
--R
--R
--R   (2)
--R  "<mrow><mrow><msup><mrow><mi>y</mi></mrow><mrow><mn>2</mn></mrow></msup></mro
--R  w><mo>+</mo><mrow><mn>2</mn><mspace width='0.3em'/><mi>x</mi><mspace width='0
--R  .3em'/><mi>y</mi></mrow><mo>+</mo><mrow><msup><mrow><mi>x</mi></mrow><mrow><m
--R  n>2</mn></mrow></msup></mrow></mrow>"
--R                                                                 Type: String
--E 2
--S 3 of 21

(x+y)**2
--R
--R
--R         2           2
--R   (3)  y  + 2x y + x
--R                                                    Type: Polynomial(Integer)
--E 3
--S 4 of 21
display(coerce(%)$MMLFORM)$MMLFORM
--R
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mrow><msup><mrow><mi>y</mi></mrow><mrow><mn>2</mn>...
--R</math>
--R                                                                   Type: Void
--E 4
)set output mathml on
--S 5 of 21
(x+y)**2
--R
--R
--R         2           2
--R   (5)  y  + 2x y + x
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mrow><msup><mrow><mi>y</mi></mrow><mrow><mn>2</mn>...
--R</math>
--R
--R                                                    Type: Polynomial(Integer)
--E 5
--S 6 of 21
integrate(x**x,x)
--R
--R
--R           x
--R         ++    %I
--R   (6)   |   %I  d%I
--R        ++
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mo>&#x0222B;</mo><mrow><msup><mrow><mi>x</mi></mrow>...
--R</math>
--R
--R                                         Type: Union(Expression(Integer),...)
--E 6
--S 7 of 21
integral(x**x,x)
--R

--R
--R           x
--R         ++    %I
--R   (7)   |   %I  d%I
--R        ++
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mo>&#x0222B;</mo><mrow><msup><mrow><mi>x</mi></mrow>...
--R</math>
--R
--R                                                    Type: Expression(Integer)
--E 7
--S 8 of 21
(5+sqrt 63 + sqrt 847)**(1/3)
--R
--R
--R         +----------+
--R        3|   +-+
--R   (8)  \|14\|7  + 5
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mroot><mrow><mrow><mrow><mrow><mrow><mn>14</mn></mrow>...
--R</math>
--R
--R                                                        Type: AlgebraicNumber
--E 8
--S 9 of 21
set [1,2,3]
--R
--R
--R   (9)  {1,2,3}
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mo>{</mo><mn>1</mn><mo>,</mo><mn>2</mn><mo>,...
--R</math>
--R
--R                                                   Type: Set(PositiveInteger)
--E 9
--S 10 of 21
multiset [x rem 5 for x in primes(2,1000)]
--R
--R
--R   (10)  {0,40: 1,47: 2,42: 3,38: 4}
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mo>{</mo><mn>0</mn><mo>,</mo><mrow><mrow><mn>40</mn></mrow>...
--R</math>
--R
--R                                                      Type: Multiset(Integer)
--E 10

--S 11 of 21
series(sin(a*x),x=0)
--R
--R
--R                3        5        7          9            11
--R               a   3    a   5    a    7     a     9      a      11      12
--R   (11)  a x - -- x  + --- x  - ---- x  + ------ x  - -------- x   + O(x  )
--R                6      120      5040      362880      39916800
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mrow><mi>a</mi><mspace width='0.3em'/><mi>x</mi></mrow>...
--R</math>
--R
--R                       Type: UnivariatePuiseuxSeries(Expression(Integer),x,0)
--E 11
--S 12 of 21
matrix [[xi+yj for i in 1..10] for j in 1..10]
--R
--R
--R   (12)
--R   [
--R     [yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi,
--R      yj + xi, yj + xi]
--R     ,
--R
--R     [yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi,
--R      yj + xi, yj + xi]
--R     ,
--R
--R     [yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi,
--R      yj + xi, yj + xi]
--R     ,
--R
--R     [yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi,
--R      yj + xi, yj + xi]
--R     ,
--R
--R     [yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi,
--R      yj + xi, yj + xi]
--R     ,
--R
--R     [yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi,
--R      yj + xi, yj + xi]
--R     ,
--R
--R     [yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi,
--R      yj + xi, yj + xi]
--R     ,
--R
--R     [yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi,

--R      yj + xi, yj + xi]
--R     ,
--R
--R     [yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi,
--R      yj + xi, yj + xi]
--R     ,
--R
--R     [yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi, yj + xi,
--R      yj + xi, yj + xi]
--R     ]
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mo>[</mo><mtable><mtr><mtd><mrow><mi>yj</mi><mo>+</mo>...
--R</math>
--R
--R                                            Type: Matrix(Polynomial(Integer))
--E 12
--S 13 of 21
y:=operator 'y
--R
--R
--R   (13)  y
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--R<mi>y</mi>
--R</math>
--R
--R                                                          Type: BasicOperator
--E 13
--S 14 of 21
D(y(x,z),[x,x,z,x])
--R
--R
--R   (14)  y        (x,z)
--R          ,1,1,2,1
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<msub><mi>y</mi><mrow><mo>,</mo><mn>1</mn><mo>,</mo><mn>1</mn><mo>,...
--R</math>
--R
--R                                                    Type: Expression(Integer)
--E 14
)clear all
--S 15 of 21
y:=operator 'y
--R
--R
--R   (1)  y
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">

--R<mi>y</mi>
--R</math>
--R
--R                                                          Type: BasicOperator
--E 15
--S 16 of 21
D(y x,x,2)
--R
--R
--R         ,,
--R   (2)  y  (x)
--R
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<msup><mi>y</mi><mrow><mo>&#x02032;</mo><mo>&#x02032;</mo>...
--R</math>
--R
--R                                                    Type: Expression(Integer)
--E 16
--S 17 of 21
x:=series 'x
--R
--R
--R   (3)  x
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--R<mi>x</mi>
--R</math>
--R
--R                       Type: UnivariatePuiseuxSeries(Expression(Integer),x,0)
--E 17
--S 18 of 21
sin(1+x)
--R
--R
--R   (4)
--R                        sin(1)  2   cos(1)  3   sin(1)  4   cos(1)  5
--R     sin(1) + cos(1)x - ------ x  - ------ x  + ------ x  + ------ x
--R                           2           6          24          120
--R   +
--R       sin(1)  6   cos(1)  7   sin(1)  8   cos(1)  9    sin(1)  10      11
--R     - ------ x  - ------ x  + ------ x  + ------ x  - ------- x   + O(x  )
--R         720        5040        40320      362880      3628800
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mrow><mo><mo>sin</mo></mo><mo>(</mo><mrow><mn>1</mn></mrow><mo>)...
--R</math>
--R
--R                       Type: UnivariatePuiseuxSeries(Expression(Integer),x,0)
--E 18

)clear all
--S 19 of 21
series(1/log(y),y=1)
--R
--R
--R   (1)
--R            - 1   1    1            1        2    19        3    3         4
--R     (y - 1)    + - - -- (y - 1) + -- (y - 1)  - --- (y - 1)  + --- (y - 1)
--R                  2   12           24            720            160
--R   +
--R        863         5    275         6    33953         7     8183         8
--R     - ----- (y - 1)  + ----- (y - 1)  - ------- (y - 1)  + ------- (y - 1)
--R       60480            24192            3628800            1036800
--R   +
--R        3250433         9            10
--R     - --------- (y - 1)  + O((y - 1)  )
--R       479001600
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mrow><msup><mrow><mrow><mo>(</mo><mi>y</mi><mo>-</mo>...
--R</math>
--R
--R                       Type: UnivariatePuiseuxSeries(Expression(Integer),y,1)
--E 19
)clear all
--S 20 of 21
y:UTS(FLOAT,'z,0):=exp(z)
--R
--R
--R   (1)
--R                    2                            3
--R     1.0 + z + 0.5 z  + 0.1666666666 6666666667 z
--R   +
--R                                4                               5
--R     0.0416666666 6666666666 7 z  + 0.0083333333 3333333333 34 z
--R   +
--R                                 6                               7
--R     0.0013888888 8888888888 89 z  + 0.0001984126 9841269841 27 z
--R   +
--R                                   8                                  9
--R     0.0000248015 8730158730 1587 z  + 0.0000027557 3192239858 90653 z
--R   +
--R                                   10      11
--R     0.2755731922 3985890653 E -6 z   + O(z  )
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mn>1.0</mn><mo>+</mo><mi>z</mi><mo>+</mo><mrow><mn>0.5</mn>...
--R</math>

--R
--R                                    Type: UnivariateTaylorSeries(Float,z,0.0)
--E 20
--S 21 of 21
c:=continuedFraction(314159/100000)
--R
--R
--R              1 |     1  |     1 |     1  |     1 |     1 |     1 |
--R   (2)  3 + +---+ + +----+ + +---+ + +----+ + +---+ + +---+ + +---+
--R            | 7     | 15     | 1     | 25     | 1     | 7     | 4
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mn>3</mn><mo>+</mo><mfrac><mn>1</mn><mrow><mn><mn>7</mn></mn>...
--R</math>
--R
--R                                             Type: ContinuedFraction(Integer)
--E 21
)spool
)lisp (bye)

———-

References^{[1] nothing}

9